This morning I found all the watchdogs had tripped during the night.
I restored them all.
I can't damp ITMX. I noticed that its driving matrix is all 1s and -1s as the the right values had been lost in some previous burtrestoring.
Rana fixed the problem. He found that the side damping was saturating. He lowered the gain a little for a while, waited for the the damping to slow down the optic and then he brought the gain back where it was.
He also upadted the MEDM screen snapshot.